OpenStudy (blackops2_143):

What is 2 1/4 • (-3 2/5)? A. -7 13/20 B. -6 13/20 C. -6 1/10 D. -5 2/20

OpenStudy (blackops2_143):

Help please?

OpenStudy (anonymous):

cross mulitply

OpenStudy (sagewilson):

Okay, fractions into decimals. 2 1/4 = 2.25, -3 2/5 = -3.40 So multiply 2.25 by -3.40.

OpenStudy (blackops2_143):

Okay thank you both

OpenStudy (blackops2_143):

I multiplyed it and got -7.65 @sagewilson

OpenStudy (sagewilson):

Well if you divide 13 by 20 you'll get .65

OpenStudy (blackops2_143):

Okay